AnakKrakatau1883:

The work of a student to solve the equation 3(2x − 4) = 8 + 2x + 6 is shown below: Step 1: 3(2x − 4) = 8 + 2x + 6 Step 2: 5x − 7 = 14 + 2x Step 3: 5x − 2x = 14 + 7 Step 4: 3x = 21 Step 5: x = 7 In which step did the student first make an error and what is the correct step? Step 2; 6x − 12 = 14 + 2x Step 2; 6x − 7 = 2(6 + x + 4) Step 3; 5x − 2x = 14 − 7 Step 3; 5x + 2x = 14 + 7

Sailor:

Ok so, in PEMDAS, you do what first?

Sailor:

Im assuming Imma have to guide you through this....

Sailor:

So, in step one, you would distribute the 3 to 2x and -4.

Sailor:

Was step one done correctly?

Sailor:

Ok... think about it this way: 3 x 2 = 6 and -4 x 3 =-12

AnakKrakatau1883:

Yes. Step one was done correctly.

Sailor:

Ok, so, if we take a look, if it's distributed properly, step 2 would look like this: 6x-12 = 8+2x+6 Then when you combine like terms, it would look like this: 6x-12=2x+14 Does step 2 look correct?

AnakKrakatau1883:

Step 2 is correct when we correctly solve Step 1.

Sailor:

But is the person's step 2 correct in the problem?

AnakKrakatau1883:

No.

Sailor:

@anakkrakatau1883 wrote:
No.
Ok, so step 2 was solved wrong, but how do we fix that?

AnakKrakatau1883:

We have to use the Order of Operations, aka PEMDAS.

Sailor:

@anakkrakatau1883 wrote:
We have to use the Order of Operations, aka PEMDAS.
Ok, so what did the person mess up on?

AnakKrakatau1883:

In the equation 3(2x − 4) = 8 + 2x + 6, the person messed up by using the distributive property.

Sailor:

@anakkrakatau1883 wrote:
In the equation 3(2x − 4) = 8 + 2x + 6, the person messed up by using the distributive property.
Ok, so that would mean that the first option for step 2 to be corrected is what?

AnakKrakatau1883:

The first option is "Step 2; 6x − 12 = 14 + 2x", which is correct in this case.

Sailor:

That's it!

AnakKrakatau1883:

Teamwork makes the dream work!

Sailor:

@anakkrakatau1883 wrote:
Teamwork makes the dream work!
I'm glad I could hep you!

Sailor:

*help

AnakKrakatau1883:

Have a blessed day!

Sailor:

@anakkrakatau1883 wrote:
Have a blessed day!
:0 ty you too